# A Non-Expert’s Introduction to Data Ethics for Mathematicians

### v2

@@ 1. Introduction
The use of digital data to examine and help understand human behavior is both powerful and dangerous [Edi21]. Every day, it seems like there is a new nightmare with problematic uses of data and algorithms. The use of predictive policing to identify criminal activity can exacerbate existing racial and ethnic inequities [Nat18], algorithmic social-credit ratings of individuals have frightening dystopian uses [Wikc], and other manifestations of “algocracy” (i.e., algorithmic government) have many human ramifications [ET19, ${}^{+}$19, Has21, SMBM23].
@@
The ever-increasing use of and reliance on “big data” and the accelerating application of tools such as machine learning and artificial intelligence (AI) to make societally consequential decisions can cause very significant harm and exacerbate societal inequities [O’N16, Ree18]. Problems like algorithmic bias and the collection, measurement, and use of enormous amounts of data about humans and their behavior have significant societal consequences [${}^{+}$21, ${}^{+}$21, SVW21, Bir21, WJ23]. New technologies like “deep fakes” [fPM20], synthetic media in which a person in an existing image or video is replaced with somebody else’s likeness, [^1] A light-hearted example is Gollum’s precious cover of the song “Nothing Compares to U” [Dor20]. have terrifying potential to cause harm. There are significant risks and potential nefarious uses of technology like generative AI and large language models (LLMs) [BGMMS21, ${}^{+}$23, Fer23]. Modern data analysis also has been accompanied by the reincarnation of pseudosciences, such as physiognomy, which can now be performed at a large scale [SH23]. Along with such dangers come enormous potential benefits, and there are nascent efforts to create documents like an “AI Bill of Rights” to help guide the design and deployment of automated systems in a way that protects people [Off22].
@@
The situation is already scary. In 2016, Microsoft released a Twitter bot that “learned” from its interactions with other Twitter accounts; in less than a day, it was regularly producing racist tweets [Vin16]. In December 2020, Stanford University’s use of an algorithm to determine which people would receive the first batches of a COVID-19 vaccine resulted in an inequitable vaccination roll-out that prioritized high-ranking doctors over frontline human-facing medical personnel [Che20]. The output of algorithms have also led to decisions to cut off medical care [RH23]. In 2021, the outgoing Editor-in-Chief of a scientific journal publicly released refereeing data, including the numbers of decisions to accept or reject papers by each of the journal’s referees during his tenure [Sch21]. In 2023, Microsoft limited its Bing AI chatbot after its unsettling conversations [Les23] and the LLM ChatGPT invented a sexual-harassment scandal and named a real professor in it [VO23]. In today’s world, it often seems necessary to laugh to keep from crying [Wei21].
@@
The mathematical, statistical, and computational sciences are interconnected with human communities and society at large [cam, eth, MC23, Mos21]. Neither our education nor our research occur in a vacuum, and many of our algorithms and other methods are now applied to social systems [${}^{+}$21, CM23]. However, it is not traditional in mathematical-science education to discuss data ethics and other ethical considerations [RSMCM22, M2̈2]. What we choose to teach (and choose not to teach) impacts what our mentees do with their education. Consequently, it is crucial for mathematical scientists (and others) to think carefully about ethics and engage with it throughout their careers. See [LMP18] for a short book on ethics and data science, [FT16] for a theme journal issue on data ethics, [V2́1] for a wide-ranging handbook of digital ethics, and [MCF22] for a recent discussion of the ethics landscape in mathematics and advocacy of a “Hippocratic Oath” for mathematics.
@@
In this chapter, I cover a diverse variety of topics. I start with data ethics and education, which is important for almost the entire mathematical-science community. In Section 2, I discuss data ethics in the mathematics community and mathematical-science education. I highlight the importance of curricular data ethics in Section 2.1, and I point to existing educational material in Section 2.2. In Section 3, I describe some of our efforts at UCLA on data ethics and society. I also briefly mention a few efforts by others and point out an important caveat in efforts in “data for social good”. In Section 4, I discuss various ethical issues and tensions in research. Some of these topics, such as research replication and appropriately acknowledging others, are directly relevant to everybody in the mathematical-science community. Other topics are most directly relevant to people who use data in their research, although it is still important for others in the mathematical sciences to have some familiarity with them. I discuss research replication and open data and code in Section 4.1, acknowledgements and licensing material for open use in Section 4.2, privacy concerns and its tension with open data and code in Section 4.3, and controversies in studies and in reactions to studies in Section 4.4. In Section 4.5, I summarize the key ideas of Section 4. In Section 5, I discuss some ethical principles and other considerations in the use of human data in scientific research. In Section 6, I highlight a few research articles about issues in data ethics. I conclude in Section 7.
@@ 2.1. The Importance of Data Ethics in the Mathematics Community and Mathematical-Science Curricula
The mathematical, statistical, and computational sciences are intertwined with human society [cam, eth, MC23]. This notion is not new. Throughout his life, educator and civil-rights activist Bob Moses championed mathematical literacy as a civil right for all public-school children, with particular advocacy for those who are most vulnerable [${}^{+}$23]. In a 2021 article that was published a month after his death [Mos21], Moses wrote
@@
It is important for mathematical scientists (and others) to think carefully about ethics and engage with it throughout their careers. We need to incorporate data ethics into the core education in the mathematical, statistical, and computational sciences. We are increasingly using social, animal, and human data (including potentially personal data) in our research. We need to think carefully about when it is appropriate to use such data and when it is not appropriate, and there needs to be systematic training to help mathematical scientists confront these issues. See [LMP18] for a short book on ethics and data science, [FT16] for a theme journal issue on data ethics, [V2́1] for a wide-ranging handbook of digital ethics, [MCF22] for a discussion of the ethics landscape in mathematics and advocacy of a “Hippocratic Oath” for mathematics (also see [D’A22]), and [Sku21] for an article with a useful discussion and pointers to several relevant resources. See [Ass18] for the Code of Ethics and Professional Conduct of the Association for Computing Machinery (ACM). A recent paper [BPT22] used the ethical-practice standards of the ACM and the American Statistical Association (ASA), which both represent disciplines that are relevant to mathematics, in a survey of mathematicians about ethical standards. To help acknowledge issues like data ethics in the mathematical sciences, it is also relevant to include pertinent information on departmental websites. For an example website, which I helped produce at University of Oxford in the aftermath of a study [TMP12] by my collaborators and me that uses Facebook data, see [MI].
@@ 3.1.1. An Undergraduate Course: Societal Impacts of Data
I helped design a new undergraduate major in “Data Theory” [^3] See [UCL] for a description of UCLA’s undergraduate major in Data Theory. at UCLA. All students in our Data Theory major must take a new upper-division course, which was designed by Mark Handcock of the Department of Statistics & Data Science, called “Societal Impacts of Data”. UCLA’s course catalog has the following description of the course:
@@ 3.3. An Important Caveat
There are a diverse variety of both research and practical efforts that fall under the auspices of “data for social good” [ACX23]. Harnessing data for social good is both important and laudable, but there are many challenges [${}^{+}$20, BBS22]. For example, it is vital to engage meaningfully and respectfully with communities, rather than attempting to help people by trying to be a “new sheriff in town”, which can be very harmful.
@@ 4.1. Replication of Research
It is crucial to be able to replicate scientific research, and improvement in current practices is necessary to produce reproducible and reliable computational science [CGH21]. Consequently, it is important to precisely explain all details of analysis, implementation, and data cleaning in scholarly works; it is also important to openly provide source code and data. The inclusion of precise explanations and source material is necessary for research dissemination. When a mathematician publishes a theorem and its proof, they also give others a license to use it freely in their own work. The sharing and portability of knowledge lies at the core of both science and mathematics. Accordingly, it is a professional obligation to share research in a usable form, including by providing source code and data. Depriving readers of this material is analogous to publishing a theorem statement without a proof or publishing a theorem without permission to use it.
@@ 4.2. Acknowledgements, Giving Credit, and Licenses to Use and Share Material
The increasing prominence of LLMs has brought new ethical concerns in acknowledgements and the creative process. [^5] More generally, the use generative AI and LLMs also brings significant ethical concerns about data ownership and copyright infringement, such as through the training data that is used for AI-generated art and text [Fra22, GM23]. Scientific journals, funding agencies, and universities are scrambling to develop policies that govern the use of LLMs and other text-generation tools in submissions [Bra23].
@@ 4.4. Controversies in Studies and in Reactions to Studies
Other research about online social networks has also been controversial. I will mention two examples that have been influential scientifically because of their research findings. One of these examples involved experimental manipulation of feeds by seeding posts on social media with a small number of fake initial upvotes or fake initial downvotes [AW12]. In this study, the researchers found that initial upvotes had a persistent effect on the overall positivity of the votes of posts, whereas the initial downvotes were overturned. The other example is the “Tastes, Ties, and Times” study of several waves of Facebook data from students at an Ivy League university in the United States [${}^{+}$08]. See [Per11] for one discussion of the data-privacy controversy of this study and its associated data set.
@@ 5.3. Institutional Review Boards (IRBs)
In many situations, it is a professional requirement to obtain permission to undertake a study in the first place. Such permission gives an ethical floor to satisfy; it is not a ceiling. It may be legally and professionally permissible to do something, but it is important to hold oneself to higher standards when it comes to whether or not it is actually the right decision to do it. For example, the limits to informed consent with human data [${}^{+}$22] may influence such a decision. Whatever you decide for your own work, make sure that you think carefully about it.
In universities in the United States, a researcher who is working with personal data needs to check with their university’s Institutional Review Board (IRB) to ensure that they are conducting research in an ethical way. Universities in other countries, private companies, government laboratories, and other organizations often have bodies that are similar to IRBs, but the procedures and especially the specific details are very different [Gri16]. A university IRB may inform you that you do not need to submit a formal application for a research project to be approved, or they may inform you that a formal application is necessary. Let your IRB know what data you have (or what data you plan to acquire and how you plan to acquire it) and what you plan to do with it. Different IRBs can rule differently. When an IRB grants permission to undertake a study, they have decided that a proposed project is above the ethical floor and hence that it is permissible to do that project. One’s own standards should be higher. (Again, ethical approval is a floor, rather than a ceiling.) In this light, it is worth examining the discussion following a controversial IRB-approved study of “emotional manipulation” in which researchers adjusted user feeds on Facebook [Luc14, KGH14].
@@ 6. Some Research and Lay Articles that are Relevant to Data Ethics and Data Privacy
There is a wealth of research about data ethics and related topics in computer science, sociology, and other fields [JIV19, ${}^{+}$21, ${}^{+}$21, ${}^{+}$20]. There is no way that I can possibly be exhaustive, [^8] I selected some of the highlighted articles, including both research studies and lay articles, from the website for Johan Ugander’s course on data ethics and data privacy [Uga20]. so I will highlight a selection of studies to give a taste of existing research. I will also briefly discuss a few articles in blogs, news websites, and other nontechnical venues. Unsurprisingly, much of this research is simultaneously fascinating and concerning [Exp18].
@@ 6.1. Algorithmic Biases and Other Biases
There is much research on the mitigation algorithmic bias, such as in the algorithmic hiring of people for jobs [RBKL20] and in classifying individuals (e.g., for admission to a university) [${}^{+}$12]. To encourage transparent reporting, clarify intended use cases, and minimize use in inappropriate contexts, some researchers have proposed the inclusion of “model cards” to accompany trained machine-learning models [${}^{+}$19]. Such model cards are short documents that give intended use cases; benchmarked evaluations across cultural, demographic, phenotypic, and intersectional groups; and other salient information.
Biases in computer systems, machine learning, and AI go far beyond only algorithmic bias [FN96]. See [BHN23, CDNSG23] for reviews of “fairness” in machine learning, and see [${}^{+}$20] for an introductory survey of bias in data-driven AI systems. As advocated in [KA21], it is important to go beyond notions of mere algorithmic “fairness” (which focuses on intra-group versus inter-group differences). One must also analyze (1) inequality and the causal impact of algorithms and (2) the distribution of power. It is also important to be cognizant of traps that can beset work on “fair” machine learning in sociotechnical systems [${}^{+}$19]. In “fair” machine learning, it is necessary to carefully examine context-specific consequences [CDNSG23].
@@ 6.2. Human Privacy, Personal Characteristics, and Personalization
Other salient research focuses on human privacy. As has been studied thoroughly [EN16], people are tracked extensively when they visit websites. It is possible to infer private traits and attributes from digital records of human behavior [KSG13]. One can also steal the identities of visitors to websites [Nar10], and trackers can use social-network structure to de-anonymize data from browsing the World Wide Web [SSGN17]. Companies can exploit the information that they observe (or infer) from website visits. For example, about a decade ago, the online travel agent Orbitz showed higher prices for flights to users of Macintosh computers than to users of other types of computers [Whi12]. It is imperative that individuals and other stakeholders have agency in how their personal data are used [SVW21], although this can come into tension with the scholarly desire to promote open data (see Section 4). For example, the sharing of data from the African continent has often been driven by non-African stakeholders [${}^{+}$21]. Additionally, conventional studies of algorithmic fairness are centered on Western culture; data proxies are different in different cultures, and it is important to consider local context when building data models [${}^{+}$21]. See [${}^{+}$21, ${}^{+}$21] for discussions of access, ethics, and best practices in the algorithmic measurement of human data.
@@
There are many ways to infer the characteristics of people and communities, as well as social and other ties between people, using data analysis. For example, it is possible to infer social ties by examining the geographical proximity of individuals in time and space using offline or online data [${}^{+}$10]. Additionally, researchers have used machine learning and Google Street View to estimate the demographic composition of neighborhoods across the United States [${}^{+}$17]. Unsurprisingly, there are significant gaps and biases (and associated exclusion issues) in such geographic data [GD22]. Moreover, as I discussed in Section 4, the fact that individuals appear in multiple social networks (e.g., multiple social-media platforms and multiple databases) gives considerable ability to identify them or information about them even when they appear to be “anonymous” in those individual networks [NS08]. For example, the inference of the sexual orientation of some individuals by combining Netflix data — which was released as part of a public competition to improve Netflix’s algorithmic ability to infer user ratings of movies — of movie rentals with movie rating data from the Internet Movie Database (IMDB) led to legal troubles, and Netflix ultimately cancelled a planned sequel competition [Wikb]. More recently, researchers have successfully inferred gender using mobile-payment data [SM23]. Moreover, one can use interactions to identify people even across long time periods [${}^{+}$22]. Tools like data analysis and machine learning can yield crucial and actionable insights (e.g., of racial disparities in police stops [${}^{+}$20]), but they need to be applied in a careful and respectful way.
@@
A related issue is personalization [Uga17], such as in the targeted advertisements that inundate people on social-media platforms. These advertisements can inform us of desirable products (such as t-shirts, dice, or plushies in my case), but the online personalization of commercial and political advertising can cause serious problems. Infamously, Donald Trump’s 2016 presidential campaign employed the company Cambridge Analytica for targeted online political advertising. Additionally, the movie Straight Outta Compton was advertised differently to people based on their inferred demographic characteristics [McA16]. There is much pertinent research. For example, there are studies of the effectiveness of spam-based marketing [${}^{+}$09], the detection of spam in social-bookmarking websites (such as Pinterest and Digg) [MCM09], and other aspects of spam and related online marketing. Users on social-media platforms like Facebook have personalized feeds, and a study that involved the adjustment of such feeds [KGH14] led to a major controversy, including with the issue of manipulating people’s emotions [Luc14].

### v3

@@ 1. Introduction
The use of digital data to examine and help understand human behavior is both powerful and dangerous [Edi21]. Every day, it seems like there is a new nightmare with problematic uses of data and algorithms. The use of predictive policing to identify criminal activity can exacerbate existing racial and ethnic inequities [Nat18], algorithmic social-credit ratings of individuals have frightening dystopian uses [Wikc], and other manifestations of “algocracy” (i.e., algorithmic government) have many human ramifications [ET19, OBS+19, Has21, SMBM23].
@@
The ever-increasing use of and reliance on “big data” and the accelerating application of tools such as machine learning and artificial intelligence (AI) to make societally consequential decisions can cause very significant harm and exacerbate societal inequities [O’N16, Ree18]. Problems like algorithmic bias and the collection, measurement, and use of enormous amounts of data about humans and their behavior have significant societal consequences [WSAO+21, LHF+21, SVW21, Bir21, WJ23]. New technologies like “deep fakes” [fPM20], synthetic media in which a person in an existing image or video is replaced with somebody else’s likeness, [^1] A light-hearted example is Gollum’s precious cover of the song “Nothing Compares to U” [Dor20]. have terrifying potential to cause harm. There are significant risks and potential nefarious uses of technology like generative AI and large language models (LLMs) [BGMMS21, EMFG+23, Fer24]. Modern data analysis also has been accompanied by the reincarnation of pseudosciences, such as physiognomy, which can now be performed at a large scale [SH22]. Along with such dangers come enormous potential benefits, and there are nascent efforts to create documents like an “AI Bill of Rights” to help guide the design and deployment of automated systems in a way that protects people [Off22].
@@
The situation is already scary. In 2016, Microsoft released a Twitter bot that “learned” from its interactions with other Twitter accounts; in less than a day, it was regularly producing racist tweets [Vin16]. In December 2020, Stanford University’s use of an algorithm to determine which people would receive the first batches of a COVID-19 vaccine resulted in an inequitable vaccination roll-out that prioritized high-ranking doctors over frontline human-facing medical personnel [Che20]. The output of algorithms have also led to decisions to cut off medical care [RH23]. In 2021, the outgoing Editor-in-Chief of a scientific journal publicly released refereeing data, including the numbers of decisions to accept or reject papers by each of the journal’s referees during his tenure [Sch21]. In 2023, the LLM ChatGPT invented [^2] It seems that LLMs and generative AI tools often simply make things up. a sexual-harassment scandal and named a real professor in it [VO23] and Microsoft limited its Bing AI chatbot after its unsettling conversations [Les23]. In today’s world, it often seems necessary to laugh to keep from crying [Wei21].
@@
The mathematical, statistical, and computational sciences are interconnected with human communities and society at large [cam, eth, MC23, Mos21]. Neither our education nor our research occur in a vacuum, and many of our algorithms and other methods are now applied to social systems [WSAO+21, CM23]. However, it is not traditional in mathematical-science education to discuss data ethics and other ethical considerations [RSMCM22, M2̈2]. What we choose to teach (and choose not to teach) impacts what our mentees do with their education. Consequently, it is crucial for mathematical scientists (and others) to think carefully about ethics and engage with it throughout their careers [CM24]. See [LMP18] for a short book on ethics and data science, [FT16] for a theme journal issue on data ethics, [V2́1] for a wide-ranging handbook of digital ethics, and [MCF22] for a recent discussion of the ethics landscape in mathematics and advocacy of a “Hippocratic Oath” for mathematics.
@@
In this chapter, I cover a diverse variety of topics. I start with data ethics and education, which is important for almost the entire mathematical-science community. In Section 2, I discuss data ethics in the mathematics community and mathematical-science education. I highlight the importance of curricular data ethics in Section 2.1, and I point to existing educational material in Section 2.2. In Section 3, I describe some of our efforts at University of California, Los Angeles (UCLA) on data ethics and society. I also briefly mention a few efforts by others and point out an important caveat in efforts in “data for social good”. In Section 4, I discuss various ethical issues and tensions in research. Some of these topics, such as research replication and appropriately acknowledging others, are directly relevant to everybody in the mathematical-science community. Other topics are most directly relevant to people who use data in their research, although it is still important for others in the mathematical sciences to have some familiarity with them. I discuss research replication and open data and code in Section 4.1, acknowledgements and licensing material for open use in Section 4.2, privacy concerns and its tension with open data and code in Section 4.3, and controversies in studies and in reactions to studies in Section 4.4. In Section 4.5, I summarize the key ideas of Section 4. In Section 5, I discuss some ethical principles and other considerations in the use of human data in scientific research. In Section 6, I highlight a few research articles about issues in data ethics. I conclude in Section 7.
@@ 2.1. The Importance of Data Ethics in the Mathematics Community and Mathematical-Science Curricula
The mathematical, statistical, and computational sciences are intertwined with human society [cam, eth, MC23]. This notion is not new. Throughout his life, educator and civil-rights activist Bob Moses championed mathematical literacy as a civil right for all public-school children, with particular advocacy for those who are most vulnerable [MWW+23]. In a 2021 article [Mos21] that was published a month after his death, Moses wrote
@@
It is important for mathematical scientists (and others) to think carefully about ethics and engage with it throughout their careers [CM24]. We need to incorporate data ethics into the core education in the mathematical, statistical, and computational sciences. We are increasingly using social, animal, and human data (including potentially personal data) in our research. We need to think carefully about when it is appropriate to use such data and when it is not appropriate, and there needs to be systematic training to help mathematical scientists confront these issues. See [LMP18] for a short book on ethics and data science, [FT16] for a theme journal issue on data ethics, [V2́1] for a wide-ranging handbook of digital ethics, [MCF22] for a discussion of the ethics landscape in mathematics and advocacy of a “Hippocratic Oath” for mathematics (also see [D’A22]), and [Sku21] for an article with a useful discussion and pointers to several relevant resources. See [Ass18] for the Code of Ethics and Professional Conduct of the Association for Computing Machinery (ACM). Buell et al. [BPT22] leveraged the ethical-practice standards of the ACM and the American Statistical Association (ASA), which both represent disciplines that are relevant to mathematics, in a survey of mathematicians about ethical standards. See [TPB24a, TPB24b] for further discussions of the guidelines that they developed for ethical mathematical practice, and see [WT24] for associated teaching materials. To help acknowledge issues like data ethics in the mathematical sciences, it is also relevant to include pertinent information on departmental websites. For an example website, which I helped produce at University of Oxford in the aftermath of a study [TMP12] by my collaborators and me that uses Facebook data, see [MI].
@@ 3.1.1. An Undergraduate Course: Societal Impacts of Data
I helped design a new undergraduate major in “Data Theory” [^4] See [Unia] for a description of UCLA’s undergraduate major in Data Theory. at UCLA. All students in our Data Theory major must take a new upper-division course, which was designed by Mark Handcock of the Department of Statistics & Data Science, called “Societal Impacts of Data”. UCLA’s course catalog has the following description of the course:
@@ 3.3. An Important Caveat
There are a diverse variety of both research and practical efforts that fall under the auspices of “data for social good” [ACX23]. Harnessing data for social good is both important and laudable, but there are many challenges [ABK+20, BBS22]. For example, it is vital to engage meaningfully and respectfully with communities, rather than attempting to help people by trying to be a “new sheriff in town”, which can be very harmful.
@@ 4.1. Replication of Research
It is crucial to be able to replicate scientific research, and improvement in current practices is necessary to produce reproducible and reliable computational science [CGH21]. Obviously, it is necessary to be honest about data and other aspects of research, but professional obligations go far beyond mere honesty. It is important to precisely explain all details of analysis, implementation, and data cleaning in scholarly works; it is also important to openly provide source code and data. See [ATG24] for a set of recommendations for sharing code. The inclusion of precise explanations and source material is necessary for research dissemination. When a mathematician publishes a theorem and its proof, they also give other people a license to use it freely in their own work. The sharing and portability of knowledge lies at the core of both science and mathematics. Accordingly, it is a professional obligation to share research in a usable form, including by providing source code and data. Depriving readers of such material is analogous to publishing a theorem statement without a proof or publishing a theorem without permission to use it.
@@ 4.2. Acknowledgements, Giving Credit, and Licenses to Use and Share Material
The increasing prominence of LLMs has brought new ethical concerns in acknowledgements and the creative process. [^6] The use generative AI and LLMs also brings significant ethical concerns about data ownership and copyright infringement, such as through the training data that is used for AI-generated art and text [Fra22, GM23]. It also can lead to comically disturbing situations, as illustrated by the recent discovery of child sexual-abuse material in a prominent data set (which has been downloaded by many researchers) in the AI community [Col23] and by the AI-generated image of a rat with a gigantic penis in a (subsequently retracted) published scientific paper [Pea24]. Scientific journals, funding agencies, professors, and universities are scrambling to develop policies that govern the use of LLMs and other text-generation tools in submissions [Bra23]. See [Con] for the LLM and AI policy that the Consortium for Mathematics and its Applications (COMAP) has started using for the Mathematical Contest in Modeling (MCM) and their other contests. For UCLA’s guidelines for the ethical and “safe” use of AI, see UCLA’s new AI website [Unib], which also discusses various AI resources. [^7] Amidst the very serious concerns about AI, it is also important not to lose sight of its immense promise. For example, see the curated list [RT24] (and the announcement of it at [Tao24]) of resources for AI in mathematics (e.g., to use AI to assist in mathematical reasoning).
@@ 4.4. Controversies in Studies and in Reactions to Studies
Other research about online social networks has also been controversial. I will mention two examples that have been influential scientifically because of their research findings. One of these examples involved experimental manipulation of feeds by seeding posts on social media with a small number of fake initial upvotes or fake initial downvotes [AW12]. In this study, the researchers found that initial upvotes had a persistent effect on the overall positivity of the votes of posts, whereas the initial downvotes were overturned. The other example is the “Tastes, Ties, and Times” study of several waves of Facebook data from students at an Ivy League university in the United States [LKG+08]. See [Per11] for one discussion of the data-privacy controversy of this study and its associated data set.
@@ 5.3. Institutional Review Boards (IRBs)
In many situations, it is a professional requirement to obtain permission to undertake a study in the first place. Such permission gives an ethical floor to satisfy; it is not a ceiling. It may be legally and professionally permissible to do something, but it is important to hold oneself to higher standards when it comes to whether or not it is actually the right decision to do it. For example, the limits to informed consent with human data [LAH+22] may influence such a decision. Whatever you decide for your own work, make sure that you think carefully about it.
In universities in the United States, a researcher who is working with personal data needs to check with their university’s IRB to ensure that they are conducting research in an ethical way. Universities in other countries, private companies, government laboratories, and other organizations often have bodies that are similar to IRBs, but the procedures and especially the specific details are very different [Gri16]. A university IRB may inform you that you do not need to submit a formal application for a research project to be approved, or they may inform you that a formal application is necessary. Let your IRB know what data you have (or what data you plan to acquire and how you plan to acquire it) and what you plan to do with it. Different IRBs can rule differently. When an IRB grants permission to undertake a study, they have decided that a proposed project is above the ethical floor and hence that it is permissible to do that project. One’s own standards should be higher. (Again, ethical approval is a floor, rather than a ceiling.) In this light, it is worth examining the discussion following a controversial IRB-approved study of “emotional manipulation” in which researchers adjusted user feeds on Facebook [Luc14, KGH14].
@@ 6. Some Research and Lay Articles that are Relevant to Data Ethics and Data Privacy
There is a wealth of research about data ethics and related topics in computer science, sociology, and other fields [JIV19, LHF+21, WSAO+21, ABK+20]. There is no way that I can possibly be exhaustive, [^10] I selected some of the highlighted articles, including both research studies and lay articles, from the website for Johan Ugander’s course on data ethics and data privacy [Uga20]. so I will highlight a selection of studies to give a taste of existing research. I will also briefly discuss a few articles in blogs, news websites, and other nontechnical venues. Unsurprisingly, much of this research is simultaneously fascinating and concerning [Exp18].
@@ 6.1. Algorithmic Biases and Other Biases
There is much research on the mitigation algorithmic bias, such as in the algorithmic hiring of people for jobs [RBKL20] and in classifying individuals (e.g., for admission to a university) [DHP+12]. To encourage transparent reporting, clarify intended use cases, and minimize use in inappropriate contexts, some researchers have proposed the inclusion of “model cards” to accompany trained machine-learning models [MWZ+19]. Such model cards are short documents that give intended use cases; benchmarked evaluations across cultural, demographic, phenotypic, and intersectional groups; and other salient information.
Biases in computer systems, machine learning, and AI go far beyond only algorithmic bias [FN96]. See [MMS+21, BHN23, CDNSG23] for reviews of bias and “fairness” in machine learning, and see [N+20] for an introductory survey of bias in data-driven AI systems. As advocated in [KA21], it is important to go beyond notions of mere algorithmic “fairness” (which focuses on intra-group versus inter-group differences). One must also analyze (1) inequality and the causal impact of algorithms and (2) the distribution of power. It is also important to be cognizant of traps that can beset work on “fair” machine learning in sociotechnical systems [SBF+19]. In “fair” machine learning, it is necessary to carefully examine context-specific consequences [CDNSG23].
@@ 6.2. Human Privacy, Personal Characteristics, and Personalization
Other salient research focuses on human privacy. As has been studied thoroughly [EN16], people are tracked extensively when they visit websites. User profiles, which encode both characteristics and behavior, are generated through interactions with websites and other digital systems. User profiling is ubiquitous in daily life, and it is also a vast area of research [PBL24]. It is possible to infer private traits and attributes from digital records of human behavior [KSG13]. One can also steal the identities of visitors to websites [Nar10], and trackers can use social-network structure to de-anonymize data from browsing the World Wide Web [SSGN17]. Companies can exploit the information that they observe (or infer) from website visits. For example, about a decade ago, the online travel agent Orbitz showed higher prices for flights to users of Macintosh computers than to users of other types of computers [Whi12]. It is imperative that individuals and other stakeholders have agency in how their personal data are used [SVW21], although this can come into tension with the scholarly desire to promote open data (see Section 4). For example, the sharing of data from the continent of Africa has often been driven by non-African stakeholders [AAB+21]. Additionally, conventional studies of algorithmic fairness are centered on Western culture; data proxies are different in different cultures, and it is important to consider local context when building data models [SAH+21]. See [LHF+21, WSAO+21] for discussions of access, ethics, and best practices in the algorithmic measurement of human data.
@@
There are many ways to infer the characteristics of people and communities, as well as social and other ties between people, using data analysis. For example, it is possible to infer social ties by examining the geographical proximity of individuals in time and space using offline or online data [CBC+10]. Additionally, researchers have used machine learning and Google Street View to estimate the demographic composition of neighborhoods across the United States [GKW+17]. Unsurprisingly, there are significant gaps and biases (and associated exclusion issues) in such geographic data [GD22]. Moreover, as I discussed in Section 4, the fact that individuals appear in multiple social networks (e.g., multiple social-media platforms and multiple databases) gives considerable ability to identify them or information about them even when they appear to be “anonymous” in those individual networks [NS08]. For example, the inference of the sexual orientation of some individuals by combining Netflix data — which was released as part of a public competition to improve Netflix’s algorithmic ability to infer user ratings of movies — of movie rentals with movie rating data from the Internet Movie Database (IMDB) led to legal troubles, and Netflix ultimately cancelled a planned sequel competition [Wikb]. More recently, researchers have successfully inferred gender using mobile-payment data [SM23]. Moreover, one can use interactions to identify people even across long time periods [CMM+22]. Tools like data analysis and machine learning can yield crucial and actionable insights (e.g., of racial disparities in police stops [PSO+20]), but they need to be applied in a careful and respectful way.
@@
A related issue is personalization [Uga17], such as in the targeted advertisements that inundate people on social-media platforms. These advertisements can inform us of desirable products (such as t-shirts, dice, or plushies in my case), but the online personalization of commercial and political advertising can cause serious problems. Infamously, Donald Trump’s 2016 presidential campaign employed the company Cambridge Analytica for targeted online political advertising. Additionally, the movie Straight Outta Compton was advertised differently to people based on their inferred demographic characteristics [McA16]. There is much pertinent research. For example, there are studies of the effectiveness of spam-based marketing [KKL+09], the detection of spam in social-bookmarking websites (such as Pinterest and Digg) [MCM09], and other aspects of spam and related online marketing. Users on social-media platforms like Facebook have personalized feeds, and a study that involved the adjustment of such feeds [KGH14] led to a major controversy, including with the issue of manipulating people’s emotions [Luc14].